OpenStudy (anonymous):

(x+2)/(3x+3)times(x^2+4x+3)/(x-1)

OpenStudy (anonymous):

First do (x^2+4x+3)/(x+1) ------------------------ Factor x^2+4x+3, and see if x+1 is a factor

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf \frac{(x+2)}{(3x+3)} \times \frac{(x^{2}+4x+3)}{(x-1)}= }\) \(\LARGE\color{blue}{ \bf \frac{(x+2)(x^{2}+4x+3)}{(3x+3)(x-1)} = }\) \(\LARGE\color{blue}{ \bf \frac{(x+2)(x+1)(x+3)}{3(x+1)(x-1)} }\) Do you see what terms can you cancel out ?

OpenStudy (anonymous):

yeah all the x's can cancel out

OpenStudy (anonymous):

thank you soo much this helped a lot!!!

OpenStudy (anonymous):

Well no: The (x+1) cancel since (x+1)/(x+1) = 1

OpenStudy (solomonzelman):

(x+1) cancels out, and nothing else.

OpenStudy (anonymous):

ohh okay